OpenStudy (anonymous):

Can someone help me on this please.The length of a retangle is 3 yd more than twice its width, the area of the retangle is 54yrd ^2. What are the deminisions of the retangle

OpenStudy (anonymous):

NOTE : negative value of x should be rejected

OpenStudy (anonymous):

ok so to find the width & the length I went (12-w) w=24 12w-w^2=24 so far this is what I have. Am I in the right direction then?

OpenStudy (anonymous):

(3+2w)*w=54

OpenStudy (anonymous):

U should solve that

OpenStudy (anonymous):

yeah I can get it from here. I had 36 instead of 54.

OpenStudy (anonymous):

(3+2w)*w=54 2w^2+3w-54=0

OpenStudy (anonymous):

2w^2+12w-9w-54=0 2w(w+6)-9(w+6)=0 (2w-9)(w+6)=0

OpenStudy (anonymous):

Now, can u?

OpenStudy (anonymous):

I am a little confused. Now that it is factored out, how can I get the length?

OpenStudy (anonymous):

either 2w-9=0 w=4.5 yard Or w+6=0 w=-6 {since length cannot be megative it is rejected}

OpenStudy (anonymous):

negative*

OpenStudy (anonymous):

ok. the negative threw me off. tk u for helping me on this.

OpenStudy (anonymous):

So, width=4.5 yard and length =(3+2*4.5)yard